FEL Upgrade Project Weekly Brief
March 12 - 16, 2001


The agenda and organization for the March 27-28 semiannual review for the FEL Upgrade project was completed and accepted.


We prepared for next week’s visit by Adm. Nanos and Mathis  (on March 22) and continued our organization of material for the March 27-28 semiannual review. All the team members are organizing material for presentation at the review with the goal of presenting our progress as clearly and concisely as possible. We are anxious to preserve the good humor of our hard working reviewers.

We reviewed the rebaselined cost performance plan for Phase 1 of the project and the proposed plan for the Phase 2 project with our ONR contract monitor (John Albertine).  Proposed revisions were discussed and accepted. The revised version will be presented at the March 27 review.

We were visited on Wednesday March 14 by a 6 member technical team from Raytheon who were interested in proposing a plan for laser development for the Navy.

WBS 3 (Beam Physics):

A "final" solution for the Rev 1.1 UV design was established, completing the initial accelerator physics design phase for the machine given our present resources. Activities now turn to documentation, performance analysis/component specification, machine modeling (gotta put that MSO training to work!) and disaster management ("What if CSR DOES matter?").

Machine studies time was used to resurrect difference orbit measurements and apply them to a carefully established machine diagnostic configuration.

WBS 4 (Injector):

(4.1)The assembly drawings for the gun chamber were completed and distributed for review.  Gun HVPS - The spare HVPS was crated and shipped to Glassman this week.  They will add a second multiplier stack and drive unit, test the new assembly, and ship it back to us.

WBS 5 (SRF):

We received the results of the gradient measurements on the second 7-cell test cavity sent to KEK for electropolish tests.  Like the first 7-cell cavity test, this cavity also shows 20MV/m gradients after electropolishing.

WBS 6 (RF):

Zone 3 - The contract for the Cathode Power Supply is being awarded this week to Hippotronics.  No other progress was made due to parts that have not yet been received.

Zone 4 - The test software is up and running with minor corrections to be made.  The hardware punch is being worked.  With the exception of positions 7 & 8, the external coefficients have been determined and they
are operating properly.  Position 8 has a circulator out for repair.  Position 7 coefficients are still being determined.

Injector HVPS - PSC is now asking us to block out the power supply before they will bid.  We thought they were the experts.

Injector 100 kW Klystron - CPI continues to have delivery problems with parts for these klystrons.  The Factory Acceptance Tests for the 1st unit have been delayed to 6/20 from  4/11.  They now plan to ship the last unit 8/15.

WBS 8 (I&C):

The design has begun for a new four channel VME delay/timing card. This will generate the required signals scope triggers, sample/holds/ BPMs, OBPMs, .... This unit will have sub-microsecond accuracy, while a commercial unit capable of sub-nanosecond resolution will be purchased.  These two boards will be able to cover any timing needs of I&C or users. This board will be close to the final timing card required for the new drive laser pulse controller (DLPC). These will located in each VME crate and driven from by fiber from the existing DLPC. All timing and
delays will be set from EPICS.

Work on the two LSS boxes is still progressing.  We are in the process of checking connections and should be able to apply power early next week. These will be installed in the two remaining user labs; #2 and #5.  Both of these will be commissioned and certified prior to the July run.  The issue of running multiple lasers will also be addressed.

All parts have been ordered for the 8th user power supply box.  The original chassis is in the process of being documented and we should have the first set of prints early next week. Final design for the hutch
patch panel and user boxes is complete and will go to EECAD early next week.

Much effort is going to specification of the trim power supplies needed for the additional magnets. A reasonable count and regulation specification should be ready by next week.

The parts for the Ion Pump Power Supplies continue to arrive for the production of 40 new chassis.  Quotes for assembly are being pursued.

Drawing collection continues.  We have received 8 additional drawings from Document Control this week.  Drawings for racks FL07B04 and FL07B05 will be to EECAD tomorrow.

WBS 9 (Transport):

The iron (1006 Steel) ordering document for the main dipoles was written and given to the procurement folks.
Injector Dipoles (DU/DV) DULY
Research is working on obtaining the required field integral in their magnetic model. When their new
800 MHz, large memory computer is available to handle the increased number of nodes, they will re look at the field flatness issue.  They are finishing up the drawing set and the drafting model of the GV (Small injector dipole with attention to the small details like lifting eyes and field clamp adjusters.
Optical Chicane Dipole (DW)
We continued back check of the remaining parts, finishing the coils.  The package should be ready in a week.
Arc Dipoles (GY, GX, GQ)
AES finished designing GY (180 degree dipole) coil's lead area and are now cleaning up the small details like lifting eyes and reference surfaces.

AES found in their magnetic model that thickening the field clamps to 3/4 inch reduced them below magnetic field saturation.  They are now investigating field integral issues noting that the field integral
increases too rapidly with radius.  The slots for the GG coils may be the culprit.  The slots may have to increase in width with radius.

QX (3.125" Quad)
Measurement Probe: The probe was used.  It is our best quadrupole probe yet, canceling out 95% of the dipole and quadrupole components.  The data is in analysis.

Design Package: Some work was accomplished on the QX procurement packages statement of work.  Jeff Karn is back devoting full time to the Upgrade.

We found that the magnetic model of the sextupole indicates the magnet is well within the 1% requirement at 2/3 radius.  The reason for the improvement in analysis is the new scrip (suggested by Steve Lassiter of
Physics) that starts analysis with the vector field potential rather than the field.

No progress in sending design of the Octupole to an Engineering Services Contractor.

Starting a new weekly meeting, Butch Dillon-Townes further clarified girder configurations by having all interested parties come together.

We continued work on defining the engineering services contractor task order that will generate the design of the dipole vacuum chambers.

WBS 11 (Optics):

We have checked the vacuum vessel for the mirror test stand against the released drawings, and found only two out-of-spec dimensions, both can be tolerated.  We have scheduled the installation of additional
utilities in User Lab 2 to begin next week, this will be used to support prototype cavity mirror tests.  We continue to work on concepts for installing 4 mirrors in the optical cavity assemblies.  We met with
staff from Survey and Alignment to determine the placement of fiducials on the mirror mounts.  This will facilitate proper installation and setting of the cavity length.

We have opened discussions with a second vendor for the upgrade drive laser.  Draft specifications for an RFQ are being written so they can be approved and passed on to Procurement.

Optics supported operations for J. Boyce (x-ray studies), I&C (O-BPM detector tests) and D. Douglas (difference orbit studies).  We had an arc lamp for the drive laser fail prematurely, the debris was cleaned up and a new lamp installed without additional problems.


We completed the winter user run last Friday. Operations this week were concerned with difference orbit measurements (see WBS 3) and O-BPM. A walk through of the building and user labs was done for the safety inspection on Friday. The web upgrade is coming along nicely.  We are working on getting system drawings (or thumbnails) to the page for use.  These will include the drawing number if more details are desired. Movie Gallery also includes an area/link for press releases as well as the most recent video shot in the labs.

Privacy and Security Notice

Privacy and Security Notice